# Deep Discriminative to Kernel Density Graph for In- and Out-of-distribution Calibrated Inference

### v7

@@ 4.1 Gaussian Kernel Parameter Estimation
We fit Gaussian kernel parameters to the samples that end up in the $r$-th polytope. We set the kernel center along a dimension $d$ as:
@@
where $x^{d}_{i}$ is the value of $\mathbf{x}_{i}$ along the $d$-th dimension. We set the kernel variance along a dimension $d$ as:
@@ 4.2 Sample Size Ratio Estimation
$\hat{\left(\frac{n_{ry}}{n_{y}}\right)}$ $=\frac{\tilde{w}_{ry}}{\tilde{w}_{y}}=\frac{\tilde{w}_{ry}}{\sum_{r\in\mathcal{P}}\tilde{w}_{ry}}$
$=\frac{\sum_{s\in\mathcal{P}}\sum_{i=1}^{n}w_{rs}\mathbbm{1}(\mathbf{x}_{i}\in Q_{s})\mathbbm{1}(y_{i}=y)}{\sum_{r\in\mathcal{P}}\sum_{s\in\mathcal{P}}\sum_{i=1}^{n}w_{rs}\mathbbm{1}(\mathbf{x}_{i}\in Q_{s})\mathbbm{1}(y_{i}=y)}.$ (11)
@@
For simplicity, we will describe the estimation procedure for $w_{rs}$ in the next sections. Note that if we satisfy Condition 12, then we have $\frac{\tilde{w}_{ry}}{\tilde{w}_{y}}\to\frac{n_{ry}}{n_{y}}$ as $n\to\infty$. Therefore, we modify Equation 5 as:
$\tilde{f}_{y}(\mathbf{x})=\frac{1}{\tilde{w}_{y}}\sum_{r\in\mathcal{P}}\tilde{w}_{ry}\mathcal{G}(\mathbf{x};\hat{\mu}_{r},\hat{\Sigma}_{r})\mathbbm{1}(r=\hat{r}^{*}_{\mathbf{x}}),$ (13)
where $\hat{r}^{*}_{\mathbf{x}}=\operatornamewithlimits{argmin}_{r}\|\hat{\mu}_{r}-\mathbf{x}\|$. Now we use $\tilde{f}_{y}(\mathbf{x})$ estimated using (13) in Equation (6), (7) and (8), respectively. Below, we describe how we estimate $w_{rs}$ for KDF and KDN .
Figure 1: Simulation datasets, Classification error, Hellinger distance from true posteriors, mean max confidence or posterior for A. five two-dimensional and B. a high dimensional (Trunk) simulation experiments. The simulation experiments were repeated $45$ times. The median performance is shown as a dark curve with shaded region as error bars showing the $25$-th and the $75$-th percentile. KDF and KDN improve both in- and out-of-distribution calibration of their respective parent algorithms while maintaining nearly similar classification accuracy on the simulation datasets. For Trunk simulation sample size was fixed at $5000$ for gradually higher number of dimensions and OOD test samples were sampled uniformly from a hyper-sphere with radius fixed at $20$. Trunk simulation dataset is visualized for the first two dimensions.
4.3 Kernel Density Forest
Consider $T$ number of decision trees in a random forest trained on $n$ i.i.d training samples $\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$. Each tree $t$ partitions the feature space into $p_{t}$ polytopes resulting in a set of polytopes: $\{\{Q_{t,r}\}_{r=1}^{p_{t}}\}_{t=1}^{T}$. The intersection of these polytopes gives a new set of polytopes $\{Q_{r}\}_{r=1}^{R}$ for the forest. For any point $\mathbf{x}_{r}\in Q_{r}$, we push every other sample point $\mathbf{x}_{s}\in Q_{s}$ down the trees. We define the kernel $\mathcal{K}(r,s)$ as:
$\mathcal{K}(r,s)=\frac{t_{rs}}{T},$ (14)
where $t_{rs}$ is the total number of trees $\mathbf{x}_{r}$ and $\mathbf{x}_{s}$ end up in the same leaf node. Note that $0\leq\mathcal{K}(r,s)\leq 1$. If the two samples end up in the same leaves in all the trees, i.e., $\mathcal{K}(r,s)=1$, they belong to the same polytope, i.e. $r=s$.
In short, $\mathcal{K}(r,s)$ is the fraction of total trees where the two samples follow the same path from the root to a leaf node. We exponentiate $\mathcal{K}(r,s)$ so that Condition 12 is satisfied:
$w_{rs}=\mathcal{K}(r,s)^{k\log n}.$ (15)
@@ 4.3 Kernel Density Forest
We choose $k$ using grid search on a hold-out dataset.
4.4 Kernel Density Network
Consider a fully connected $L$ layer ReLU-net trained on $n$ $iid$ training samples $\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$. We have the set of all nodes denoted by $\mathcal{N}_{l}$ at a particular layer $l$. We can randomly pick a node $n_{l}\in\mathcal{N}_{l}$ at each layer $l$, and construct a sequence of nodes starting at the input layer and ending at the output layer which we call an activation path: $m=\{n_{l}\in\mathcal{N}_{l}\}_{l=1}^{L}$. Note that there are $N=\Pi_{i=1}^{L}{|\mathcal{N}_{l}|}$ possible activation paths for a sample in the ReLU-net. We index each path by a unique identifier number $z\in\mathbb{N}$ and construct a sequence of activation paths as: =$\mathcal{M}=\{m_{z}\}_{z=1,\cdots,N}$. Therefore, $\mathcal{M}$ contains all possible activation pathways from the input to the output of the network.
While pushing a training sample $\mathbf{x}_{i}$ through the network, we define the activation from a ReLU unit at any node as ‘$1$’ when it has non-negative input and ‘00’ otherwise. Therefore, the activation indicates on which side of the affine function at each node the sample falls. The activation for all nodes in an activation path $m_{z}$ for a particular sample creates an activation mode $a_{z}\in\{0,1\}^{L}$. If we evaluate the activation mode for all activation paths in $\mathcal{M}$ while pushing a sample through the network, we get a sequence of activation modes: $\mathcal{A}_{r}=\{a_{z}^{r}\}_{z=1}^{N}$. Here $r$ is the index of the polytope where the sample falls in.
If the two sequences of activation modes for two different training samples are identical, they belong to the same polytope. In other words, if $\mathcal{A}_{r}=\mathcal{A}_{s}$, then $Q_{r}=Q_{s}$. This statement holds because the above samples will lie on the same side of the affine function at each node in different layers of the network. Now, we define the kernel $\mathcal{K}(r,s)$ as:
$\mathcal{K}(r,s)=\frac{\sum_{z=1}^{N}\mathbbm{1}(a_{z}^{r}=a_{z}^{s})}{N}.$ (16)
Note that $0\leq\mathcal{K}(r,s)\leq 1$. In short, $\mathcal{K}(r,s)$ is the fraction of total activation paths which are identically activated for two samples in two different polytopes $r$ and $s$. We exponentiate the kernel using Equation 15.
Pseudocodes outlining the two algorithms are provided in Appendix C.
4.5 Geodesic Distance
Consider $\mathcal{P}_{n}=\{Q_{1},Q_{2},\cdots,Q_{R}\}$ as a partition of $\mathbb{R}^{d}$ given by a random forest or a ReLU-net after being trained on $n$ training samples. We measure distance between two points $\mathbf{x}\in Q_{r},\mathbf{x}^{\prime}\in Q_{s}$ using the kernel introduced in Equation 14 and Equation 16 [22] and call it ‘Geodesic’ distance:
$d(r,s)=-\mathcal{K}(r,s)+\frac{1}{2}(\mathcal{K}(r,r)+\mathcal{K}(s,s))=1-\mathcal{K}(r,s)$ (17)
Proposition 4.1.
$(\mathcal{P}_{n},d)$ is a metric space.
Proof 4.2.
See Appendix A.1 for the proof.
As $d$ cannot distinguish between points within the same polytope, it has a resolution similar to the size of the polytope. For discriminating between two points within the same polytope, we fit a Gaussian kernel within the polytope (described above). As $h_{n}\to 0$, the resolution for Geodesic distance improves. In Section 5, we will empirically show that using Geodesic distance scales better with higher dimension compared to that of Euclidean distance.
Lemma 4.3.
For any two points, $\mathbf{x}_{1}\in Q_{r}$ and $\mathbf{x}_{2}\in Q_{s}$, if $d(r,s)=0$, that is, $\mathcal{K}(r,s)=1$, we have $r=s$.
Given $n$ independent and identically distributed training samples $\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$, we define the distance of an inference point $\mathbf{x}$ from the training points as: $d_{\mathbf{x}}=\min_{i=1,\cdots,n}\|\mathbf{x}-\mathbf{x}_{i}\|$.
Proposition 4.4 (Asymptotic OOD Convergence).
Given non-zero and bounded bandwidth of the Gaussians, then we have almost sure convergence for $\hat{g}_{y}$ as:
@@ 5 Empirical Results
We conduct several experiments on simulated, OpenML-CC18 [23]  [^1] https://www.openml.org/s/99 and vision benchmark datasets to gain insights on the finite sample performance of KDF and KDN. The details of the simulation datasets and hyperparameters used for all the experiments are provided in Appendix B. For Trunk simulation dataset, we follow the simulation setup proposed by Trunk [24] which was designed to demonstrate ‘curse of dimensionality’. In the Trunk simulation, a binary class dataset is used where each class is sampled from a Gaussian distribution with higher dimensions having increasingly less discriminative information. We use both Euclidean and Geodesic distance in (4) on simulation datasets and use only Geodesic distance for benchmark datasets. For the simulation setups, we use classification error, hellinger distance [25, 26] from the true class conditional posteriors and mean max confidence or posterior [4] as performance statistics. While measuring in-distribution calibration for the datasets in OpenML-CC18 data suite, as we do not know the true distribution, we used maximum calibration error as defined by Guo et al. [18] with a fixed bin number of $R=15$ across all the datasets. Given $n$ OOD samples, we define OOD calibration error (OCE) to measure OOD performance for the benchmark datasets as:
$\text{OCE}=\frac{1}{n}\sum_{i=1}^{n}\left|\max_{y\in\mathcal{Y}}(\hat{P}_{Y|X}(y|\mathbf{x}_{i}))-\max_{y\in\mathcal{Y}}(\hat{P}_{Y}(y))\right|.$ (18)
For the tabular and the vision datasets, we have used ID calibration approaches, such as Isotonic [15, 16] and Sigmoid [17] regression, as baselines. Additionally, for the vision benchmark dataset, we provide results with OOD calibration approaches such as: ACET [4], ODIN [6], OE (outlier exposure) [10]. We report traditionally used OOD detection statistics such as area under curve (AUROC) and false positive rate at $95\%$ precision (FPR@95). For each approach, $90\%$ of the training data was used to fit the model and the rest of the data was used to calibrate the model.
@@ 5.1.1 Simulation Study
Figure 1 leftmost column shows $10000$ training samples with $5000$ samples per class sampled within the region $[-1,1]\times[-1,1]$ from the six simulation setups described in Appendix B. Therefore, the empty annular region between $[-1,1]\times[-1,1]$ and $[-2,2]\times[-2,2]$ is the low density or OOD region in Figure 1. Figure 1 quantifies the performance of the algorithms which are visually represented in Appendix Figure 4. KDF and KDN maintain similar classification accuracy to those of their parent algorithms. We measure hellinger distance from the true distribution for increasing training sample size within $[-1,1]\times[-1,1]$ region as a statistics for in-distribution calibration. Column $3$ of left and Column $2$ of right block in Figure 1 show KDF and KDN are better at estimating the high density region of training distribution compared to their parent methods. In all of the simulations, using geodesic distance measure results in better performance compared to those while using Euclidean distance. For measuring OOD performance, we keep the training sample size fixed at $1000$ and normalize the training data by the maximum of their $l2$ norm so that the training data is confined within a unit circle. For inference, we sample $1000$ inference points uniformly from a circle where the circles have increasing radius and plot mean max posterior for increasing distance from the origin. Therefore, for distance up to $1$ we have in-distribution samples and distances farther than $1$ can be considered as OOD region. As shown in Column $3$ of Figure 1, mean max posteriors or confidence for KDF and KDN converge to the maximum of the class priors, i.e., $0.5$ as we go farther away from the training data origin.
Row $6$ of Figure 1 shows KDF-Geodesic and KDN-Geodesic scale better with higher dimensions compared to their Euclidean counterpart algorithms respectively.
@@ 5.1.2 OpenML-CC18 Data Study
Figure 2: Performance summary of KDF and KDN on OpenML-CC18 data suite. The dark curve in the middle shows the median of performance on $45$ datasets. The shaded region shows the error bar consisting of the $25$-th and the $75$-th percentile of the performance statistics. We conducted one-sided Wilcoxon rank-sum test using improvements in different datasets at the final sample size as test statistics. Each panel shows a p-value for each baseline approach which correspond to testing whether baseline model performs worse than our approach. Left: KDF and KDN maintains or improves performance of their parent algorithms for classification. Middle: KDF and KDN significantly improve the in-distribution calibration for random forest and ReLU-net. Right: Both of the proposed approaches yield highly calibrated confidence in the OOD region.
We use OpenML-CC18 data suite for tabular benchmark dataset study. We exclude any dataset which contains categorical features or NaN values  [^2] We also excluded the dataset with dataset id $23517$ as we could not achieve better than chance accuracy using RF and DN on that dataset. and conduct our experiments on $45$ datasets with varying dimensions and sample sizes. For the OOD experiments, we follow a similar setup as that of the simulation data. We normalize the training data by their maximum $l_{2}$ norm and sample $1000$ testing samples uniformly from each hypersphere where the hyperspheres have increasing radius starting from $1$ to $5$. For each dataset, we measure improvement with respect to the parent algorithm as:
@@
where $\mathcal{E}_{p}=$classification error, MCE or OCE for the parent algorithm and $\mathcal{E}_{M}$ represents the performance of the approach in consideration. Note that positive improvement implies the corresponding approach performs better than the parent approach. We report the median of improvement on different datasets along with the error bar in Figure 2. The extended results for each dataset is shown separately in the appendix. Figure 2 left column shows on average KDF and KDN has nearly similar or better classification accuracy to their respective parent algorithm whereas Isotonic and Sigmoid regression have lower classification accuracy most of the cases. However, according to Figure 2 middle column, KDF and KDN have similar in-distribution calibration performance to the other baseline approaches. Most interestingly, Figure 2 right column shows that KDN and KDF improves OOD calibration of their respective parent algorithms by a huge margin while the baseline approaches completely fails to address the OOD calibration problem.

### v9

@@ 2.7 Gaussian Kernel Parameter Estimation
We fit Gaussian kernel parameters to the samples that end up in the $r$-th polytope. We set the kernel center along the $d$-th dimension:
@@
where $x^{d}_{i}$ is the value of $\mathbf{x}_{i}$ along the $d$-th dimension. We set the kernel variance along the $d$-th dimension:
@@ 2.8 Sample Size Ratio Estimation
$\frac{\tilde{n}_{ry}}{\tilde{n}_{y}}=\frac{\sum_{s\in\mathcal{P}}\sum_{i=1}^{n}w_{rs}\mathbbm{1}(\mathbf{x}_{i}\in Q_{s})\mathbbm{1}(y_{i}=y)}{\sum_{r\in\mathcal{P}}\sum_{s\in\mathcal{P}}\sum_{i=1}^{n}w_{rs}\mathbbm{1}(\mathbf{x}_{i}\in Q_{s})\mathbbm{1}(y_{i}=y)}.$ (16)
@@
Note that if we satisfy the conditions in Equation 17, then we have $\frac{\tilde{n}_{ry}}{\tilde{n}_{y}}\to\frac{n_{ry}}{n_{y}}$ as $n\to\infty$. We exponentiate $\mathcal{K}(r,s)$ so that the conditions in Equation 17 is satisfied:
$w_{rs}=\mathcal{K}(\boldsymbol{x}_{i},\boldsymbol{x}_{j})^{\gamma\log n}.$ (18)
@@
We choose $\gamma$ using grid search on a hold-out dataset. Therefore, we modify Equation 7 as:
$\hat{f}_{y}(\mathbf{x})=\frac{1}{\tilde{n}_{y}}\sum_{r\in\mathcal{P}}\tilde{n}_{ry}\mathcal{G}(\mathbf{x};\hat{\mu}_{r},\hat{\Sigma}_{r})\mathbbm{1}(r=\hat{r}^{*}_{\mathbf{x}}),$ (19)
where $\hat{r}^{*}_{\mathbf{x}}=\operatornamewithlimits{argmin}_{r}\|\hat{\mu}_{r}-\mathbf{x}\|$. Now we use $\hat{f}_{y}(\mathbf{x})$ estimated using (19) in Equation (8), (9) and (10), respectively.
[…]
Proposition 2 (Asymptotic OOD Convergence).
Given non-zero and bounded bandwidth of the Gaussians, then we have:
@@
Proof.
See Appendix B.2 for the proof. ∎
Note that the above proposition guarantees asymptotic performance of the proposed approach. In contrast, typical existing approaches lack any such guarantees. Below we conduct empirical experiments including simulations and benchmark datasets to gain insights about the non-asymptotic performance of the proposed approach.
Algorithm 2 Fit a KDX model.
1:
4:$\mathcal{G}$triangleright a KDX model
5:function KDX.fit($\theta,\mathbf{X},\mathbf{y}$)
7:         for $j=1,\ldots,n$ do
8:              $k_{ij}\leftarrow$ computeKernel($\mathbf{x}_{i},\mathbf{x}_{j},\theta$)
9:         end for
10:     end for
11:
12:     $\{\{k_{rs}\}_{r=1}^{\tilde{p}}\}_{s=1}^{\tilde{p}}\leftarrow$ getPolytopes($\mathbf{k}$) $\triangleright$ Identify the polytopes populated by the training data by merging two samples with $k_{ij}=1$
13:
14:     for $r=1,\ldots,\tilde{p}$ do $\triangleright$ Iterate over each polytope
15:         $\mathcal{G}.\hat{\mu}_{r},\mathcal{G}.\hat{\Sigma}_{r},\mathcal{G}.\hat{n}_{ry}\leftarrow$ estimateParameters($\mathbf{X},y,\{k_{rs}\}_{s=1}^{\tilde{p}}$) $\triangleright$ Estimate model parameters
16:     end for
17:     return $\mathcal{G}$
18:end function
2.9 Performance Measure
For the simulation setups, we evaluate performance with classification error, Hellinger distance (Kailath, 1967; Rao, 1995) between the true and estimated class conditional posteriors, and mean max confidence (MMC) (Hein et al., 2019). We use expected calibration error (ECE) to evaluate in-distribution calibration for the OpenML-CC18 data suite, with $R=20$ bins for all datasets Guo et al. (2017). Given $n$ OOD samples $\{\mathbf{x}_{i}\}_{i=1}^{n}$, we define OOD calibration error (OCE) to measure OOD performance for the benchmark datasets as:
$\text{OCE}=\frac{1}{n}\sum_{i=1}^{n}\left|\max_{y\in\mathcal{Y}}(\hat{P}_{Y|X}(y|\mathbf{x}_{i}))-\max_{y\in\mathcal{Y}}(P_{Y}(y))\right|.$ (20)
Note that one needs to know the true prior to calculating OCE. For tabular data experiments, training points are sampled based on empirical class priors in the overall dataset. In vision experiments, we assume equal class priors and sample training points accordingly.
@@ 3 Empirical Results
For empirical experiments, we first demonstrate that our methods satisfy the theoretical guarantees on simulated datasets where we know the ground truth, and then we apply our approach on benchmark datasets such as OpenML-CC18 (Bischl et al., 2017)  [^1] https://www.openml.org/s/99 and different vision datasets. The details of the simulation data sets and hyperparameters used for all experiments are provided in Appendix D. For tabular and vision data sets, we have used ID calibration approaches, such as Isotonic Zadrozny & Elkan (2001); Caruana (2004) and Sigmoid regression Platt et al. (1999), as baselines. Additionally, for the vision benchmark dataset, we provide results with OOD calibration approaches such as: ACET Hein et al. (2019), ODIN Liang et al. (2017), OE (outlier exposure) Hendrycks et al. (2018). For each approach, $70\%$ of the training data was used to fit the model and the rest of the data was used to calibrate the model.
@@ 3.1 Empirical Study on Tabular Data
Figure 3: KDF and KDN improve both in-distribution and out-of-distribution calibration of the parent model while preserving classification accuracy. We evaluate performance using simulation datasets, reporting classification error, Hellinger distance from true posteriors, and mean max confidence (MMC). Results are presented for: A. Two-dimensional simulations. In the fourth and seventh columns, MMC is measured as a function of increasing distance from the origin and B. High-dimensional (Trunk) simulations — visualized along the first two dimensions. OOD points are randomly sampled from a sphere of radius 20, and MMC is measured across increasing dimensions. For Hellinger distance, lower values indicate a closer match to the true training distribution, while in OOD regions, a value near 0.5 signifies better OOD calibration. The median performance of $10$ repetitions is shown as a dark curve with shaded region as error bars.
@@ 3.1.1 Simulation Study
For simulation study on tabular data, we use $6$ simulation datasets. Three of the simulations are visualized in Figure 3A and see Appendix D for additional simulations and details. We sample $10$,$000$ training samples with half of the samples from each class. The samples are sampled within the region $[-1,1]\times[-1,1]$, rendering the empty annular region between $[-1,1]\times[-1,1]$ and $[-2,2]\times[-2,2]$ as the low density or OOD region. In the Trunk simulation (Figure 3 B), each of two classes is sampled from a Gaussian distribution, with higher dimensions having increasingly less discriminative information (Trunk, 1979). We use both Euclidean and Geodesic distances to detect the nearest polytope center. To measure in-distribution performance, we report classification error and the Hellinger distance between the estimated and true distributions. To measure OOD performance, we normalize the training data to the maximum of their norm $\ell_{2}$. For inference, we sample $1000$ inference points uniformly from a circle and compute the mean max posterior for increasing distance from the origin. Therefore, for a distance of up to $1$ we have in-distribution samples and distances further than $1$ are OOD. For the Trunk simulation, we randomly sampled OOD points from a sphere of fixed radius $20$ and measured mean max confidence for increasing dimensions.
KDF and KDN achieve classification accuracy comparable to their parent algorithms (Figure 3A). However, they provide a more accurate estimation of the in-distribution (ID) region. The use of geodesic distance further enhances performance relative to Euclidean distance. As data points move farther from the training set, the mean maximum confidence for KDF and KDN approaches the maximum of the class priors, reaching $0.5$. Furthermore, KDF-Geodesic and KDN-Geodesic exhibit better scalability in higher dimensional spaces compared to their Euclidean counterparts (Figure 3 B).
Hereafter, we only use Geodesic distance to find the nearest polytope in our algorithms.
@@ 3.1.2 OpenML-CC18 Data Study
Figure 4: Performance summary of KDF and KDN on OpenML-CC18 data suite. The dark curve in the middle shows the median of performance on $45$ datasets with the shaded region as error bar.
Figure 5: KDN filters out inference points with different kinds of shifts from the training data. Simulated images: (A) circle with radius $10$, (B) rectangle with sides $(20,50)$ and (C) out-of-distribution test points, ellipse with minor and major axis $(10,30)$. Mean max confidence of KDN are plotted for semantic shift of the inference points created by (D) changing the color intensity, (E) taking convex combination of circle and rectangle, i.e., pixel-wise convex combination of the images in A and B, (F) changing one of the axes of the ellipse.
We use OpenML-CC18 data suite for tabular benchmark dataset study. We exclude any dataset which contains categorical features or NaN values  [^2] We also excluded the dataset with dataset id $23517$ as we could not achieve better than chance accuracy using RF and DN on that dataset. and conduct our experiments on $45$ datasets with varying dimensions and sample sizes. We repeat the experiment for each dataset $10$ times. For the OOD experiments, we follow a similar setup as that of the simulation data. We normalize the training data by their maximum $\ell_{2}$ norm and sample $1000$ testing samples uniformly from hyperspheres where each hypersphere has increasing radius starting from $1$ to $5$. For each dataset, we measure improvement with respect to the parent algorithm:
@@
where $\mathcal{E}_{p}$ represents either classification error, ECE or OCE for the parent algorithm and $\mathcal{E}_{M}$ represents the performance of the approach in consideration. Positive improvement implies that the corresponding approach performs better than the parent approach. We report the median of improvement on different datasets along with the interquartile ranges (Figure 4, extended results for each dataset in the appendix). On average KDF and KDN have similar or better classification accuracy compared to their respective parent algorithms, whereas Isotonic and Sigmoid regression have lower classification accuracy most of the cases. KDF and KDN have similar ID calibration performance to the other baseline approaches, all better than the parent algorithm for all sample sizes. Most interestingly, KDN and KDF improve OOD calibration of their respective parent algorithms by a huge margin while the baseline approaches completely fail in OOD calibration.